OpenStudy (anonymous):

if 8n + 22 = 70 find the value of 3n

OpenStudy (jack1):

@ayeissel first find the value of n then multiply by 3 to find the value of 3n if: 8n + 22 = 70 what happens if you take 22 from each side of the equals sign?

OpenStudy (08surya):

calculate value of n and then multiply it with 3

HanAkoSolo (jamierox4ev3r):

So first you would solve this like a regular equation: 8n+22=70 -22 -22 --------- 8n=48 if 8n =48, then n=? @ayeissel ?